What is the total impedance of a circuit with a 20 ohm resistor, a 50 ohm capacitive reactance, 30 ohm inductive reactance in parallel?
1
Expert's answer
2020-08-25T11:13:11-0400

For the parallel circuit: "\frac{1}{Z}=\sqrt{\frac{1}{R}^2+\frac{1}{X}^2}" , where Z is impedance, R - resistance, X - reactance. "\frac{1}{X}=\frac{1}{X_L}-\frac{1}{X_C}." Hence, "\frac{1}{Z}=\sqrt{\frac{1}{R}^2+(\frac{1}{X_L}-\frac{1}{X_C})^2}=\sqrt{\frac{1}{20}^2+(\frac{1}{30}-\frac{1}{50})^2}="

"=\sqrt{(\frac{1}{400}+(\frac{20}{1500})^2}=0.052 Ohm^{-1}". Thus, the total impedance is "Z=\frac{1}{0.052}=19.23Ohm".